# Test-Time Training Scaling for Chemical Exploration in Drug Design

### v1

@@ 4.2 Test-time training with RL for chemical exploration
We trained a population of independent agents using the best-performing baseline (ACEGENMolOpt) configuration. In this case, each agent had a budget of 10,000 molecules. The results of different population sizes up to 128 agents is shown in Figure 2. We found that the benchmark performance continued to increase log-linearly from $\sim 1.6$ to $\sim 3.5$ up until the maximum population size was tested, almost reaching the maximum of score of 4. Interestingly, when the population size reached 32, the estimated maximum score for this RL configuration was exceeded. This is indicative of the power of leveraging the variance between different agents to overcome the more ambiguous reward signal, which did not turn out to be a limiting factor. Figure 15 shows the average similarity to each target during training across multiple agents, indicating the variance observed between multiple agents with different agents maximizing similarity to different molecular targets, resulting in overall increased exploration and increased benchmark scores. This is exemplified in Figure 3 which shows that a population of 128 agents identifies all 3 A2A targets. Interestingly, for all population sizes in the AP task, all agents maximized the same molecular target, suggesting a particular difficulty in optimizing the other.
Figure 2: Benchmark performance on MolExp as the number of independent agents increases. The score reported is the product of the maximum similarity to each target molecule summed over the tasks. Note that performance increase scales log-linearly with increasing independent agents. Estimated max is calculated based on a single agent optimizing each task’s molecular target individually.
Figure 3: Example molecules generated during the A2A task from the MolExp benchmark. In columns, the molecule with the maximum similarity to each target is shown, alongside which agent sampled it. 1 agent almost solves target 3, 4 agents solve target 3 and almost target 2, meanwhile 128 agents solve all 3 targets.
Increasing the number of agents, each with an individual budget, effectively increases the total budget. Therefore, we tested the single agent with the cumulative budget of a population of agents, for comparison. Figure 4 shows that a single agent trained for the equivalent budget does not increase performance similarly. In this case, scaling the duration of training for a single agent results in a saturation of benchmark performance after approximately 40,000 molecules evaluated, as agents expectedly fail to continue exploring once a target has been maximized (Figure 16). This is only improved slightly if state-of-the-art exploration strategies are employed such as RND, the most performant exploration strategy (Svensson et al., 2024; Thomas et al., 2025), which provides an exploration bonus to the reward. Where Figure 17 shows only small fluctuations in similarity due to RND, before subsequent maximization again to the same ’chosen’ molecular target. This is instead of maximization of similarity to a given molecular target and then a switch from one to another, as intended with exploration methods.
Figure 4: Performance comparison between using multiple independent agents or running one single agent for longer. Note that running multiple independent agents with a budget of 10,000 each outperforms running a single agent with a combined budget of $10,000\times N_{agents}$. Additional exploration strategies such as RND don’t improve performance to the level of multiple independent agents.
4.3 Cooperative multi-agent RL for chemical exploration
@@
Cooperation within a population of RL agents can theoretically improve performance by leveraging useful information shared between agents, such as, states already visited and their corresponding reward (Oroojlooy & Hajinezhad, 2023). We explored several strategies to leverage cooperative learning to improve the efficiency of chemical exploration. Considering the maximum number of task objectives in the benchmark was 4, we investigated strategies utilizing 4 RL agents. Where, the best-case scenario is if each agent learns a different objective.
The MolExp score of the different cooperative strategies are shown in Figure 5. All of the strategies tested were similar or worse in performance, with no statistically significant improvement over independent agents. The best strategies were POPNORM, ENTS, and DIFFN. Where lower $\lambda$ values performed best for ENTS showing divergent behavior for a single replicate in Figure 18 with little restriction in learning capability, although this is observed at larger values of $\lambda$. Meanwhile either the lowest (0.001) or highest (1) tested $\lambda$ value for DIFFN performed well. Interestingly, there didn’t seem to be any obvious correlation between the $\lambda$ value and restriction on learning capability for the values tested (Figure 20). Note that the best values here differ from the default of 0.01 used by Hu et al. (2024), also used with 4 CLM agents.
The Noise strategy exhibited higher standard deviation in benchmark score. Both Purge and Shared replay buffer strategies dropped performance, with Shared being among the worst performing benchmarks as anticipated. Where in the example replicate shown in Figure 18 every agent optimizes the same molecular target and hence converged in chemical space as anticipated, with the novelty bonus failing to ameliorate this effect. Purge may not offer strict enough segregation of replay buffers to diverge agent behavior, as unique but similar molecules may still exist in different replay buffers. On the other hand, strategies like RND and DvD that should more strongly encourage divergent behavior appear to ultimately restrict learning within the defined budget (Figure 18), without resulting in the desired outcome of different agents optimizing different molecular targets more than independent agents. Lastly, no tested values of $\lambda$ for CES or DIFFS outperformed independent agents. Where higher values also led to restricted learning capability, especially in the case of DIFFS as shown in Figure 19 and Figure 20.
@@
Figure 5: Performance comparison of different multi-agent cooperative strategies on the MolExp benchmark. Each plot is a different cooperative strategy using 4 agents. The dashed line represents the average performance of 4 independent agents. Note that no strategy tested significantly outperformed a population of 4 independent agents as measured by a Bonferroni-corrected, one-tailed t-test.
@@ 5 Conclusion
In this paper, we present the challenge of exploring chemical space with a CLM to identify all possible candidate molecules in the context of drug discovery. We introduced a new benchmark to measure chemical exploration highlighting the poor intrinsic exploration of current baseline algorithms. We showed that this benchmark could be practically solved by a population size of up to 128 independent agents which performed superior to a single agent trained for longer, even with exploration bonuses. Lastly, the cooperative strategies we proposed and tested were unable to significantly outperform independent agents in our setting.
The scaling of populations of agents in our setting has a direct relation to the scaling of TTT in LLMs. In our case, a CLM agent is pre-trained to form a generalist model that can sample molecules in a similar chemical space to the training dataset. We temporarily update the CLM agent with RL on a specific test-time task (each of the four benchmark tasks), which can be viewed as iterations of training and sampling, and collect the outcomes (molecules sampled) but throw away the updated parameters. However, as opposed to scaling just the inference-time as typically done, we scaled the whole TTT process by increasing the number of agents undergoing TTT on a specific test-time task. Interestingly, observing the same log-linear scaling law in performance with the number of samples.
@@
Overall, this work takes an important step towards the future scenario of identifying all possible molecules of interest in drug-like chemical space. Our MolExp benchmark is the first to explicitly ask this question, which we hope will guide algorithmic improvement until chemical space can be efficiently and thoroughly explored for a given context. Populations of RL agents are a promising solution towards this future.
Impact statement
This paper presents work whose goal is to advance the field of Machine Learning in application to drug discovery. We don’t anticipate any direct societal consequences of our work that must be specifically highlighted here.
Appendix A Chemical language model implementation
We trained a gated recurrent unit (GRU) network on molecules within the ChEMBL34 database (Gaulton et al., 2012) to act as the CLM.
@@ Appendix A Chemical language model implementation
First, we filtered the database to form the training dataset and ensured no molecules existed with more 10 rotatable bonds, a logP above 5.5, a molecular weight outside of 150-650 Da, or contained atoms not in a comment set $a\in{C,N,S,O,F,Cl,Br,H}$. All molecules were neutralized if possible, converted to non-isomeric SMILES (i.e., containing no stereoinformation) and de-duplicated. This resulted in a dataset of 1,711,022 unique molecules. Finally, the target molecules of the MolExp benchmark Appendix C were purposefully confirmed to be present within the training dataset, to ensure that objectives were theoretically achievable.
@@
The GRU consisted of an embedding of 256 and 3 layers of GRU cells with hidden dimension 512. This was trained with restricted SMILES randomization (Arús-Pous et al., 2019b) for 10 epochs with a batch size of 128. This was implemented such that a random SMILES string was each time a dataset molecule was sampled during training. An ADAM optimizer was employed with a learning rate of 0.001 and step scheduler that dropped the learning rate every 500 steps. The learning rate, loss and SMILES validity during training is shown in Figure 6.
@@
Figure 6: Pre-training of the CLM on a training dataset of ChEMBL34 molecules.
@@ Appendix B Cooperative strategies
As initially proposed by Hu et al. (2024), an additional loss penalty was added which seeks to maximize the difference in log-probability for on-policy states $\mathcal{S}^{k}$ between the $k$-th agent and previous agents, scaled by the reward $R(\tau)$.
@@
Taking inspiration from Parker-Holder et al. (2020) we implemented Diversity via Determinant to enforce divergent behavior of agent policies. A random selection of 100 on-policy states were sampled and used to form an embedding that encoded policy behavior. This embedding was formed by concatenating the transition probabilities for possible actions $P(s^{i}_{t+1}|a^{i}_{t},s_{t})$ for a given agent policy $\pi^{i}_{\theta}$ for each sampled state $s$. Let this form the agent embedding $\mathbf{e_{i}}$ for each agent $i$ which were then used to compute pairwise similarities via a radial basis function (RBF) kernel:
@@ Appendix C Molecular exploration benchmark
In this work, the similarity function $sim$ is implemented as the Levenshtein similarity between the canonical SMILES strings. This choice is because the CLM generates and acts in a SMILES environment. We created the Levenshtein similarity by normalizing the Levenshtein distance by the length of the reference string, clipped to ensure no negative values:
@@
Note that the SMILES strings of both $m$ $m_{t}$ are canonicalized and both used non-isomeric SMILES. Moreover,we made an alternative similarity function available using the Tanimoto similarity of ECFP fingerprints for application with non-language models, although the Levenshtein similarity will still work provided a SMILES representation is given.
@@
Figure 8: Adenosine A2A (A2A) receptor drug candidates as molecular targets. All the molecular candidates are A2A receptor antagonists that bind to the same orthosteric site but with alternative binding modes. XAC binds from N253 up through the pocket with a polar tail that extends into solvent, LUAA47070 binds via a water mediated interaction with N253, while Imaradenant binds from N253 down occupying lipophilic hotspots
@@ Appendix D Baseline algorithms
REINVENT: This baseline was implemented following the REINVENT formulation (Olivecrona et al., 2017) whereby a fixed version of the prior policy $\pi_{prior}$ (i.e., the pre-trained CLM) is used to regularize the updates of the agent policy $\pi_{agent}$ being trained. This is a form of reward shaping as shown in following Equation 20. This also introduces a hyperparameter $\sigma$ that determines the balance between $\pi_{prior}-\pi_{agent}$ difference or the reward $R(\tau)$. The hyperparameters used corresponding to the default frequently used in the literature are shown in Table 3.
@@
[…]

### v2

@@ 4.2 Test-time training scaling for chemical exploration
We tested if scaling TTT with larger population sizes of independent RL agents would increase chemical exploration and improve benchmark performance. In this case, we chose the best-performing RL configuration-ACEGENMolOpt-and assigned a budget of 10,000 molecules to each agent in the population. 2(a) shows that MolExpL performance increases log-linearly from $\sim 1.6$ with 1 agent to $\sim 3.5$ with 128 agents, almost reaching a maximum score of 4. Log-linear scaling was also confirmed using the vanilla REINFORCE RL configuration, albeit with lower benchmark scores. Interestingly, when the population size reached 32 the estimated maximum score for this RL configuration was exceeded, indicating that the additional variance between independent agents overcomes the additionally reward signal ambiguity introduced from multiple molecular targets. Figure 17 shows that by increasing agents, there is more variance across the set of molecule targets. The AP task is an exception, where all 128 independent agents all maximize similarity to the same target molecule in the set. Similar results are also observed when using fingerprint similarity oracles subsection E.2. Additionally, similar scaling was achieved with up to 32 agents on the GuacaMol benchmark (subsection F.3), at which point performance on many tasks becomes saturated.
(a) Scaling RL agents
(b) Scaling budget
(c) Molecular diversity
Figure 2: Performance on the MolExpL benchmark with scaling. (a) Scaling the number of independent ACEGENMolOpt and REINFORCE RL agents, each with a budget of 10,000. (b) Scaling the total budget allocated to a single agent, a single agent with a RND exploration bonus, and a single agent with a DF. (c) The diversity of sampled compounds as measured by sphere exclusion diversity.
Increasing the number of agents, each with a budget of 10,000, increases the total budget. Therefore, we compared the alternative approach of scaling TTT via increasing the budget for a single RL agent. However, this did not result in similar MolExpL performance gain as shown in 2(b). Single RL agent performance saturates at a budget of 40,000 as the agent expectedly fails to continue exploring once one of the target molecules has been maximized (Figure 18). Moreover, 2(c) shows that molecular diversity decreases while scaling budget. Employing RND as a state-of-the-art exploration strategy, the most performant exploration strategy [35, 19], marginally improves performance and molecular diversity. Where Figure 19 shows that RND only results in fluctuations in similarity to one molecular target, as opposed to the desired behavior of switching from one molecular target in the set to another. We additionally tested a simple diversity filter (DF) that assigns a reward of 0 to non-unique canonical SMILES. However, the DF also failed to rescue TTT scaling in this dimension. Figure 3 exemplifies how 87 agents are required to solve the A2A task and recover all target molecules, while scaling one agent results in a focus only on the third target molecule.
Figure 3: Example molecules generated compared to the set of targets in the MolExpL A2A task. The maximum similarity score and corresponding agent $k$ is labeled. Note that only with 87 agents is the task solved.
4.3 Cooperative multi-agent RL for targeted chemical exploration
@@
Cooperation within a population of RL agents can theoretically improve performance by leveraging useful information shared between agents, such as, states already visited and their corresponding reward [38]. We explored several strategies to leverage cooperative learning to improve the efficiency of chemical exploration. Considering the maximum number of molecular targets in a task set was 4, we investigated strategies utilizing 4 RL agents. Where, the best-case scenario is each agent maximizing similarity to a different molecular target in the set.
We investigated an extensive range of the different cooperative strategies on the MolExpL benchmark, including the MolRL-MGPT baseline with the same budget and number of agents, shown in 4(a). All of the strategies tested were similar or worse in performance, with the best being POPNORM, ENTS, and DIFFN. However, no strategy significantly outperformed independent agents as measured by a Bonferroni-corrected, one-tailed t-test. On the other hand, 4(b) shows that many cooperative strategies successfully resulted in corresponding increases in molecular diversity, namely RND, ENTS, CES, and DIFFS. This is usually anti-correlated with MolExpL score. This highlights the difficulty and complexity in achieving targeted exploration, where diversity appears to hinder learning, or is naive rather than targeted on on favorable regions of chemical space. These trends were even more prominent when using fingerprint similarity oracles subsection G.2.
Assessing the example behavior of different agents shows that some cooperative strategies do result in divergent behavior, but to the detriment of learning efficiency by slowing or inhibiting the learning of subsequent agents in a population (see subsection G.1).
(a) Cooperative RL MolExpL benchmark performance
(b) Cooperative RL molecular diversity
@@
Figure 4: Performance comparison of different 4-agent cooperative strategies on the MolExpL benchmark, each with a budget of 10,000. The dashed line represents the average of 4 independent agents as baseline.
4.4 Maximizing predicted A2A bioactivity
To represent a practically applicable objective function, we additionally tested TTT scaling behavior when maximizing the predicted probability of A2A bioactivity as measured by the QSAR classification model from MolScore [43]. Increasing chemical exploration during optimization should increase the probability of recovering the set of known A2A drug candidates used in the MolExp A2A task, two of which are present in the training dataset positive class. Therefore, we use the same metric to measure the maximimum similarity achieved to the A2A target molecule set, we refer to this additional experiment as MolExpBio. However, this now depends on the properties of the QSAR model and therefore, is not theoretically guaranteed.
5(a) shows that scaling the number of independent RL agents increases the benchmark score, again more than scaling the budget of a single RL agent (5(b)). In other words, scaling RL agents increases the maximum similarity to the set of target molecules, and hence the chance of rediscovering the drug candidates. This represents increased chemical exploration of the reward landscape. This task highlights the translational benefit of TTT scaling with RL agents to real-world drug design objectives.
(a) Scaling RL agents
(b) Scaling budget
(c) Molecular diversity
Figure 5: Performance on the MolExpBio task with scaling. (a) Scaling the number of independent ACEGENMolOpt and REINFORCE RL agents, each with a budget of 10,000. (b) Scaling the total budget allocated to a single agent, a single agent with a RND exploration bonus, and a single agent with a DF. (c) The diversity of sampled compounds as measured by sphere exclusion diversity.
@@ 5 Conclusion
In this paper, we probe the exploration of chemical space with a CLM to identify multiple candidate molecules in the context of drug discovery. In this effort, we introduce MolExp as a new benchmark, showing poor intrinsic exploration of current baseline algorithms. We showed that scaling a population of independent RL agents was the best approach to increasing benchmark performance, practically solving it with 128 agents. This included comparisons to scaling budgets for single RL agents and an extensive range of cooperative strategies.
Moreover, we found that performance improved log-linearly, similarly observed with test-time inference scaling of LLMs. In this regard, training the pre-trained CLM agent with RL on a specific test-time task (each of the four benchmark tasks) can be viewed as TTT. We than scale the whole TTT process by increasing the number of RL agents, resulting in similar scaling laws observed with test-time inference scaling in LLMs.
@@
Overall, this work takes an important step towards identifying all possible molecules of interest in drug-like chemical space. Our MolExp benchmark is the first to explicitly ask this question, which we hope will guide algorithmic improvement until chemical space can be efficiently and thoroughly explored for a given context. TTT scaling through populations of RL agents are a promising solution towards this future.
Limitations
The similarity functions used for the tasks are unlikely to be used in real-world drug design. In practice, predictive models of bioactivity, binding affinity, and ADMET properties are more common. However, the set of all the best possible solutions for such scoring functions (e.g., molecular docking) are currently unknowable. For this reason, MolExp(L) implements similarity functions as perfect scoring functions and tasks with known ground truths. We acknowledge this difference and strengthen the translational assumption with similar empirical results on other established benchmarks and a real-world objective in A2A bioactivity prediction.
We did not investigate the effect of model architecture and training dataset in this work, which has been thoroughly investigated elsewhere. Instead we focus on maximizing chemical exploration efficiency given the same or similar starting policy. These results are expected to translate to any starting policy.
The caveat of scaling TTT with more RL agents or increased budget is increased walltime requirements, as shown in Appendix I. This reiterates the importance of trying to use cooperative methods to improve exploration efficiency, as well as, the use of this benchmark and results as a tool for further research into targeted and efficient chemical space exploration.
Code availability
The code used to generate this data is part of the ACEGEN repository https://github.com/Acellera/acegen-open. The code and training dataset for the MolExp benchmark is part of the MolScore repository https://github.com/MorganCThomas/MolScore which is integrated into ACEGEN.
Appendix A Model pre-training
A.1 Datasets
@@
MolExp benchmark The pre-training dataset for the MolExp(L) benchmark was curated from the ChEMBL34 database [44]. First, the database was filtered to ensure no molecules existed with more 10 rotatable bonds, a logP above 5.5, a molecular weight outside of 150-650 Da, or contained atoms not in a comment set $a\in{C,N,S,O,F,Cl,Br,H}$. All molecules were neutralized if possible, converted to non-isomeric SMILES (i.e., containing no stereoinformation) and de-duplicated. This resulted in a dataset of 1,711,022 unique molecules. Finally, the target molecules of the MolExp benchmark Appendix C were purposefully confirmed to be present within the training dataset, to ensure that objectives were theoretically achievable.
GuacaMol benchmark The pre-training dataset for the GuacaMol [12] benchmark was taken directly from the original publication without any further processing. This dataset was downloaded from https://figshare.com/projects/GuacaMol/56639.
A.2 ACEGEN
@@
MolExp benchmark The CLM constituting the prior policy is a gated recurrent unit (GRU) network with an embedding of 256 and 3 layers of GRU cells with hidden dimension 512. This was trained with restricted SMILES randomization [45] for 10 epochs with a batch size of 128. This was implemented such that a random SMILES string augmentation was applied each time a dataset molecule was sampled during training. An ADAM optimizer was employed with a learning rate of 0.001 and step scheduler that dropped the learning rate every 500 steps. The learning rate, training loss, and SMILES validity during training is shown in Figure 6.
@@
Figure 6: Pre-training of the GRU CLM on the MolExp benchmark pre-training dataset curated from ChEMBL34.
GuacaMol benchmark The CLM constituting the prior policy is an long short-term memory (LSTM) network with the same hyperparameters as the model used in GuacaMol benchmark. This includes an embedding of 1024 and 3 layers of LSTM cells with hidden dimension 1024, as well as a dropout rate of 0.2. The pre-trained weights were loaded from the GuacaMol benchmark resulting in the exact same prior and initial policy.
A.3 MolRL-MGPT
The MolRL-MGPT model was re-implemented following the code provided from the original publication [36]. This CLM constituting the prior policy is a GPT-style Transformer model with 8 layers each with 8 attention heads. For a standardized comparison, the model was pre-trained on the benchmark training datasets respectively. For hardware reasons the batch size was lowered from 2048 to 512, with all other hyperparameters kept equal including training for 10 epochs. The learning rate, training loss, and SMILES validity during training is shown in x.
Figure 7: Pre-training of the MolRL-MGPT CLM on the MolExp benchmark pre-training dataset and GuacaMol benchmark pre-training dataset.
@@ Appendix B Cooperative strategies
As initially proposed by Hu et al. [36], an additional loss penalty was added which seeks to maximize the difference in log-probability for on-policy states $\mathcal{S}^{k}$ between the $k$-th agent and previous agents, scaled by the reward $R(\tau)$.
@@
Taking inspiration from Parker-Holder et al. [27], we implemented Diversity via Determinant to enforce divergent behavior of agent policies. A random selection of 100 on-policy states were sampled and used to form an embedding that encoded policy behavior. This embedding was formed by concatenating the transition probabilities for possible actions $P(s^{i}_{t+1}|a^{i}_{t},s_{t})$ for a given agent policy $\pi^{i}_{\theta}$ for each sampled state $s$. Let this form the agent embedding $\mathbf{e_{i}}$ for each agent $i$ which were then used to compute pairwise similarities via a radial basis function (RBF) kernel:
@@ Appendix C Molecular exploration benchmark
In this work, the similarity function $sim$ is implemented as the Levenshtein similarity between the canonical SMILES strings (MolExpL). This was chosen because the CLM generates and acts in a SMILES environment. The Levenshtein similarity function was created by normalizing the Levenshtein distance by the length of the reference string, clipped to ensure the output value was bound between [0, 1]:
@@
Note that the SMILES strings of both $m$ $m_{t}$ are canonicalized and both used non-isomeric SMILES. Moreover,we made an alternative similarity function available using the Tanimoto similarity of ECFP fingerprints for application with non-language models (MolExp), although the Levenshtein similarity will still work provided a SMILES representation is given.
@@
Figure 10: Adenosine A2A (A2A) receptor drug candidates as molecular targets. All the molecular candidates are A2A receptor antagonists that bind to the same orthosteric site but with alternative binding modes. XAC binds from N253 up towards the extracellular side with a polar tail that extends into solvent, LUAA47070 binds via a water mediated interaction with N253, while Imaradenant binds from N253 down towards the intrahelical bundle occupying lipophilic hotspots.
@@ Appendix D Baseline algorithms
REINVENT: This baseline was implemented following the REINVENT formulation [14] whereby a fixed version of the prior policy $\pi_{prior}$ (i.e., the pre-trained CLM) is used to regularize the updates of the agent policy $\pi_{agent}$ being trained. This is a form of reward shaping as shown in following Equation 20. This also introduces a hyperparameter $\sigma$ that determines the balance between $\pi_{prior}-\pi_{agent}$ difference or the reward $R(\tau)$. The hyperparameters used corresponding to the default frequently used in the literature are shown in Table 3.
@@
AHC: Augmented Hill-Climb [17] is a variation of REINVENT that conducts on-policy hill-climbing, i.e., on-policy molecules are ranked by reward $R(\tau)$ and only the top-$k$ molecules are retained. This shift focuses learning on high-reward molecules which improves efficiency, and helps to avoid too much regularization in low-reward scenarios. All configuration parameters are shown in Table 5.
MolRL-MGPT: MolRL-MGPT is a cooperative multi-agent RL baseline with 4 GPT-style CLM agents. This follows the REINVENT formulation with the addition of a second loss term for the $k$-th agent encouraging reward scaled deviation of the current agent to previous ones. This baseline was implemented using the code provided alongside the publication [36]. The default hyperparameters provided with the code were used, as there were insufficient details for full reproducibility of paper results.
$\mathcal{L}^{k}_{DIFF_{\mathcal{N}}}=-\sigma_{2}\cdot\sum_{i=1}^{k-1}R(\tau)\cdot\lvert\log\pi_{k}(\tau)-\log\pi_{i}(\tau)\rvert$ (21)